OpenStudy (anonymous):

An angle in standard position in the standard(x,y) coordinate plane has its vertex at the origin(0,0), and its initial side on the positive x-axis. If the measure of an angle in standard position is 2,397°, an angle of what degree measure has the same terminal side?

OpenStudy (amistre64):

divide by 360 to get a remainder

OpenStudy (anonymous):

so 2,397/360?

OpenStudy (amistre64):

2397 = 360n + R where R is the same angle between 0 and 360

OpenStudy (amistre64):

so yes; but dont take it into the decimals; just determine the remainder part

OpenStudy (amistre64):

2397 = R (mod 360)

OpenStudy (anonymous):

so i take it that solve for R to get my answer?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

solve this 1: 2397 = R (mod 360) or this 1: 2397 = 360n + R

OpenStudy (anonymous):

?

OpenStudy (amistre64):

both are the same, just different notations

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i can solve either 1 for R and still get the same answer?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

2397/360 = 6.658... ; ignore the decimal and you have "n" 2397 = 360(6) + R